OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11.

OpenStudy (anonymous):

then just equate h(x) to 11 2x + 1 = 11 solve for x

OpenStudy (anonymous):

so is it 6?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

2x+1=11. Isolate x by itself on the left side.

OpenStudy (anonymous):

To do that, subtract 1 from each side. 2x+1-1=11-1

OpenStudy (anonymous):

2x=10

OpenStudy (anonymous):

Then divide both sides by 2. What do you get?

OpenStudy (anonymous):

ummmmm.....

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

kk :3